A 90.5kg man I'd standing on a surface of negligible friction kicks forward a 45gm stone lying at his feet so that he requires a speed of 3.05m/s. What velocity does the man acquire as a result
1
Expert's answer
2021-10-04T13:34:26-0400

By the law of conservation of momentum, we have:


"p_i=p_f,""m_{stone}v_{stone}+m_{man}v_{man}=0,""v_{man}=-\dfrac{m_{stone}v_{stone}}{m_{man}}=-\dfrac{0.045\ kg\cdot3.05\ \dfrac{m}{s}}{90.5\ kg}=-1.5\cdot10^{-3}\ \dfrac{m}{s}."

The sign minus means that the man moves in the opposite direction to the motion of the stone.

Answer:

"v_{man}=1.5\cdot10^{-3}\ \dfrac{m}{s}," in the opposite direction to the motion of the stone.